OpenStudy (anonymous):

Which polyhedron has incorrect entries for its number of faces, edges, or vertices?

OpenStudy (studygurl14):

are there answer choices to choose from?

OpenStudy (anonymous):

Polyhedron Number of Faces Number of Edges Number of Vertices Polyhedron 1 4 6 4 Polyhedron 2 6 12 8 Polyhedron 3 8 10 6 Polyhedron 4 12 30 20

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

theres some equation that aids in this that i can never really recall. something like v+f = 2+e

OpenStudy (anonymous):

Okay, ill try that

OpenStudy (amistre64):

google says its good, so essentially its asking you if v+f-e = 2

OpenStudy (anonymous):

Yes. That means polyhedron 3 is incorrect. Can you double check that for me??

OpenStudy (amistre64):

Polyhedron 3 f8 e10 v6 6+8 -10 = 4 this ones is bad fer sure

OpenStudy (anonymous):

Thanks!!